# Self-Normalized Resets for Plasticity in Continual Learning

### v1

@@ B.2.4 Convergence After a Positive Initialization
$F(w_{0})$ $=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w_{0}^{T}x)-\sigma(v^{T}x))^{2}]$
$=F(0)+\frac{1}{2}\mathbb{E}_{x}[\sigma(w_{0}^{T}x)^{2}]-\mathbb{E}_{x}[\sigma(w_{0}^{T}x)\sigma(v^{T}x)]$
@@
$=F(0)+\frac{\tilde{w}_{0}^{2}L^{2}}{12}-\mathbb{E}_{x}[\sigma(w_{0}^{T}x)\sigma(v^{T}x)]$
Therefore, we define $\delta=-\frac{\tilde{w}_{0}^{2}L^{2}}{12}+\mathbb{E}_{x}[\sigma(w_{0}^{T}x)\sigma(v^{T}x)]$ and we proceed to show that $\delta\geq\frac{\tilde{w}_{0}L^{2}}{24}$. We let $z=-\frac{b_{v}}{\tilde{v}}$ so that $(z,1)$ is the $x$-intercept of the line $v^{T}x$. Then, $v^{T}x=\tilde{v}(\tilde{x}-z)$. Moreover, $z\in[0,\frac{L}{6}]$ since $b_{v}\in[-\frac{L}{6},0]$ and $\tilde{v}\geq 1$. Therefore, $\forall\tilde{x}\geq z,v^{T}x\geq 0$ and $\forall\tilde{x}<z,v^{T}x<0$. Since $b_{w_{0}}=0$ and $\tilde{w}_{0}\geq 0$, then $\forall\tilde{x}\geq 0,w_{0}^{T}x\geq 0$ and $\forall\tilde{x}<0,w_{0}^{T}x<0$. Thus,
$\sigma(w_{0}^{T}x)\sigma(v^{T}x)=\begin{cases}\tilde{w}_{0}\tilde{x}\tilde{v}(\tilde{x}-z)&\text{if $\tilde{x}\geq z$}\\ 0&\text{if $\tilde{x}<z$}\end{cases}$ (21)
@@
$\mathbb{E}_{x}[\sigma(w_{0}^{T}x)\sigma(v^{T}x)]$ $=\frac{1}{2L}\tilde{w}_{0}\tilde{v}\int_{z}^{L}y^{2}-yzdy$ by (21)
@@
$\delta$ $=-\frac{\tilde{w}_{0}^{2}L^{2}}{12}+\mathbb{E}_{x}[\sigma(w_{0}^{T}x)\sigma(v^{T}x)]$
@@
By Lemma B.6, there exists some $\delta\geq\frac{\tilde{w}_{0}L^{2}}{24}$ such that $F(w_{0})\leq F(0)-\delta$. Then (22) follows by the latter inequality and a slight modification of the proof of Theorem 5.2 (along with Lemma D.2, Lemma D.4, and Lemma D.6) of Vardi et al. (2021), to extend the result to target parameters $v$ with arbitrary magnitudes. Specifically, a simple modification of the aforementioned proofs implies that setting $\gamma=\frac{\delta^{3}}{3\cdot 12^{2}(||w_{0}||+||v||+1)^{5}c^{8}c^{\prime 2}}$ where we take $c=\max_{x}||x||=\sqrt{L^{2}+1}$ and $c^{\prime}=1$ due to $x\sim\text{ Uniform}(-L,L)\times\{1\}$ guarantees (22). Finally, (23) follows by
$F(w_{t})$ $=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w_{t}^{T}x)-\sigma(v^{T}x))^{2}]$
$\leq\mathbb{E}_{x}[\frac{1}{2}(w_{t}^{T}x-v^{T}x)^{2}]$ by $\sigma(\cdot)$ 1-Lipschitz continuous, Lemma B.1
@@
B.2.5 No Resets After a Positive Initialization
Lemma B.8.
@@ B.2.5 No Resets After a Positive Initialization
$\sup_{x\in[-L,L]\times\{1\}}w^{T}x\leq\epsilon\Rightarrow F(w)\geq F(0)-\frac{\tilde{v}L}{2}\epsilon$
@@
$F(w_{t})$ $=\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))^{2}]$
$\geq\mathbb{E}_{x}[(\sigma(w^{T}x)-\sigma(v^{T}x))^{2}\mathbb{I}_{\{v^{T}x\geq\epsilon\}}]$
$\geq\mathbb{E}_{x}[(\sigma(v^{T}x)-\epsilon)^{2}\mathbb{I}_{\{v^{T}x\geq\epsilon\}}]$ by $\sup_{x\in[-L,L]\times\{1\}}w^{T}x\leq\epsilon$
Let $z=-\frac{b_{v}}{\tilde{v}}$ such that $v^{T}(z,1)=0$. Then $v^{T}(z+\frac{\epsilon}{\tilde{v}},1)=\epsilon$ and since $\tilde{v}>0$ then $\forall\tilde{x}\geq z+\frac{\epsilon}{\tilde{v}}$ we have that $v^{T}(x,1)\geq\epsilon$. Therefore,
$\mathbb{E}_{x}[(\sigma(v^{T}x)-\epsilon)^{2}\mathbb{I}_{\{v^{T}x\geq\epsilon\}}]$ $=\mathbb{E}_{x}[(v^{T}x-\epsilon)^{2}\mathbb{I}_{\{v^{T}x\geq\epsilon\}}]$
@@
According to Lemma B.7, there exists some $\delta\geq\frac{\tilde{w_{0}}L^{2}}{24}$ such that
@@
Additionally, Lemma B.7, which utilizes a modified proof of Theorem 5.2 of Vardi et al. (2021), ensures that
@@
$\sup_{x\in[-L,L]\times\{1\}}u_{t+1}^{T}x\leq\epsilon$
Then according to Lemma B.8,
@@
B.2.6 Proof of Theorem B.2
Theorem B.3 (Restatement of Theorem B.2).
@@
In order to apply Lemmas B.5, B.7, and B.9 we verify that
@@
By Lemma B.5 and Lemma B.9 and the choice of learning rate $\eta$, we only ever reset if $\tilde{w}_{t}$ is (re)initialized such that $\tilde{w}_{t}<0$. Conversely, if we (re)initialize the parameters such that $\tilde{w}_{t}>0$ then we never reset for subsequent gradient descent updates. Additionally, since $w_{t}$ is reinitialized by an independent draw from its initial distribution, then for any particular reset count $r$ we have that
@@
By the choice of learning rate $\eta$ and Lemma B.5, if $w$ is (re)initialized such that $\tilde{w}_{t}<0$ then within $T_{\text{reset}}\leq\lceil\frac{32L^{2}F((-1,0))}{\eta\epsilon^{4}}\rceil$ gradient descent updates $w$ is reset. Moreover, the loss $F(w_{t^{\prime}})$ is at most $F((-1,0))$ for time steps $t^{\prime}$ preceding a reset of $w$ as loss is maximized at $w_{t}=(-1,0)$ over initializations and choosing learning rate $\eta$ at most $\frac{1}{L^{2}+1}$, the Lipschitz constant of $\nabla F$ (Lemma B.3), guarantees that loss never exceeds $F((-1,0))$. Thus,
@@
$\leq\sum_{t=0}^{\infty}(L^{2}+1)(1-\eta\gamma)^{t}\mathbb{E}_{w_{0}}[||w_{0}-v||^{2}|\tilde{w}_{0}>0]$ by Lemma B.7

### v2

@@ E.2.4 Convergence After a Positive Initialization
$F(w_{0})$ $=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w_{0}^{\top}x)-\sigma(v^{\top}x))^{2}]$
$=F(0)+\frac{1}{2}\mathbb{E}_{x}[\sigma(w_{0}^{\top}x)^{2}]-\mathbb{E}_{x}[\sigma(w_{0}^{\top}x)\sigma(v^{\top}x)]$
@@
$=F(0)+\frac{\tilde{w}_{0}^{2}L^{2}}{12}-\mathbb{E}_{x}[\sigma(w_{0}^{\top}x)\sigma(v^{\top}x)]$
Therefore, we define $\delta=-\frac{\tilde{w}_{0}^{2}L^{2}}{12}+\mathbb{E}_{x}[\sigma(w_{0}^{\top}x)\sigma(v^{\top}x)]$ and we proceed to show that $\delta\geq\frac{\tilde{w}_{0}L^{2}}{24}$. We let $z=-\frac{b_{v}}{\tilde{v}}$ so that $(z,1)$ is the $x$-intercept of the line $v^{\top}x$. Then, $v^{\top}x=\tilde{v}(\tilde{x}-z)$. Moreover, $z\in[0,\frac{L}{6}]$ since $b_{v}\in[-\frac{L}{6},0]$ and $\tilde{v}\geq 1$. Therefore, $\forall\tilde{x}\geq z,v^{\top}x\geq 0$ and $\forall\tilde{x}<z,v^{\top}x<0$. Since $b_{w_{0}}=0$ and $\tilde{w}_{0}\geq 0$, then $\forall\tilde{x}\geq 0,w_{0}^{\top}x\geq 0$ and $\forall\tilde{x}<0,w_{0}^{\top}x<0$. Thus,
$\sigma(w_{0}^{\top}x)\sigma(v^{\top}x)=\begin{cases}\tilde{w}_{0}\tilde{x}\tilde{v}(\tilde{x}-z)&\text{if $\tilde{x}\geq z$}\\ 0&\text{if $\tilde{x}<z$}\end{cases}$ (16)
@@
$\mathbb{E}_{x}[\sigma(w_{0}^{\top}x)\sigma(v^{\top}x)]$ $=\frac{1}{2L}\tilde{w}_{0}\tilde{v}\int_{z}^{L}y^{2}-yzdy$ by (16)
@@
$\delta$ $=-\frac{\tilde{w}_{0}^{2}L^{2}}{12}+\mathbb{E}_{x}[\sigma(w_{0}^{\top}x)\sigma(v^{\top}x)]$
@@
By Lemma E.6, there exists some $\delta\geq\frac{\tilde{w}_{0}L^{2}}{24}$ such that $F(w_{0})\leq F(0)-\delta$. Then (17) follows by the latter inequality and a slight modification of the proof of Theorem 5.2 (along with Lemma D.2, Lemma D.4, and Lemma D.6) of Vardi et al. (2021), to extend the result to target parameters $v$ with arbitrary magnitudes. Specifically, a simple modification of the aforementioned proofs implies that setting $\gamma=\frac{\delta^{3}}{3\cdot 12^{2}(||w_{0}||+||v||+1)^{5}c^{8}c^{\prime 2}}$ where we take $c=\max_{x}||x||=\sqrt{L^{2}+1}$ and $c^{\prime}=1$ due to $x\sim\text{ Uniform}(-L,L)\times\{1\}$ guarantees (17). Finally, (18) follows by
$F(w_{t})$ $=\mathbb{E}_{x}[\frac{1}{2}(\sigma(w_{t}^{\top}x)-\sigma(v^{\top}x))^{2}]$
$\leq\mathbb{E}_{x}[\frac{1}{2}(w_{t}^{\top}x-v^{\top}x)^{2}]$ by $\sigma(\cdot)$ 1-Lipschitz continuous, Lemma E.1
@@
E.2.5 No Resets After a Positive Initialization
Lemma E.8.
@@ E.2.5 No Resets After a Positive Initialization
$\sup_{x\in[-L,L]\times\{1\}}w^{\top}x\leq\epsilon\Rightarrow F(w)\geq F(0)-\frac{\tilde{v}L}{2}\epsilon$
@@
$F(w_{t})$ $=\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))^{2}]$
$\geq\mathbb{E}_{x}[(\sigma(w^{\top}x)-\sigma(v^{\top}x))^{2}\mathbb{I}_{\{v^{\top}x\geq\epsilon\}}]$
$\geq\mathbb{E}_{x}[(\sigma(v^{\top}x)-\epsilon)^{2}\mathbb{I}_{\{v^{\top}x\geq\epsilon\}}]$ by $\sup_{x\in[-L,L]\times\{1\}}w^{\top}x\leq\epsilon$
Let $z=-\frac{b_{v}}{\tilde{v}}$ such that $v^{\top}(z,1)=0$. Then $v^{\top}(z+\frac{\epsilon}{\tilde{v}},1)=\epsilon$ and since $\tilde{v}>0$ then $\forall\tilde{x}\geq z+\frac{\epsilon}{\tilde{v}}$ we have that $v^{\top}(x,1)\geq\epsilon$. Therefore,
$\mathbb{E}_{x}[(\sigma(v^{\top}x)-\epsilon)^{2}\mathbb{I}_{\{v^{\top}x\geq\epsilon\}}]$ $=\mathbb{E}_{x}[(v^{\top}x-\epsilon)^{2}\mathbb{I}_{\{v^{\top}x\geq\epsilon\}}]$
@@
According to Lemma E.7, there exists some $\delta\geq\frac{\tilde{w_{0}}L^{2}}{24}$ such that
@@
Additionally, Lemma E.7, which utilizes a modified proof of Theorem 5.2 of Vardi et al. (2021), ensures that
@@
$\sup_{x\in[-L,L]\times\{1\}}u_{t+1}^{\top}x\leq\epsilon$
Then according to Lemma E.8,
@@
E.2.6 Proof of Theorem E.2
Theorem E.3 (Restatement of Theorem E.2).
@@
In order to apply Lemmas E.5, E.7, and E.9 we verify that
@@
By Lemma E.5 and Lemma E.9 and the choice of learning rate $\eta$, we only ever reset if $\tilde{w}_{t}$ is (re)initialized such that $\tilde{w}_{t}<0$. Conversely, if we (re)initialize the parameters such that $\tilde{w}_{t}>0$ then we never reset for subsequent gradient descent updates. Additionally, since $w_{t}$ is reinitialized by an independent draw from its initial distribution, then for any particular reset count $r$ we have that
@@
By the choice of learning rate $\eta$ and Lemma E.5, if $w$ is (re)initialized such that $\tilde{w}_{t}<0$ then within $T_{\text{reset}}\leq\lceil\frac{32L^{2}F((-1,0))}{\eta\epsilon^{4}}\rceil$ gradient descent updates $w$ is reset. Moreover, the loss $F(w_{t^{\prime}})$ is at most $F((-1,0))$ for time steps $t^{\prime}$ preceding a reset of $w$ as loss is maximized at $w_{t}=(-1,0)$ over initializations and choosing learning rate $\eta$ at most $\frac{1}{L^{2}+1}$, the Lipschitz constant of $\nabla F$ (Lemma E.3), guarantees that loss never exceeds $F((-1,0))$. Thus,
@@
$\leq\sum_{t=0}^{\infty}(L^{2}+1)(1-\eta\gamma)^{t}\mathbb{E}_{w_{0}}[||w_{0}-v||^{2}|\tilde{w}_{0}>0]$ by Lemma E.7